# ATLAS: Adaptive Trading with LLM AgentS Through Dynamic Prompt Optimization and Multi-Agent Coordination

### v1

@@ 5 Results
Tables 1 and 2 compare three prompting strategies within ATLAS (fixed baseline, reflection, and Adaptive-OPRO) across multiple LLM families and regime-specific equity studies, alongside non-LLM baselines, under our proposed experimental design. Results show that Adaptive-OPRO consistently improves upon fixed prompts across models and market conditions, while reflection often deteriorates performance or provides inconsistent value. Non-LLM strategies demonstrate regime-dependent performance, with different technical approaches succeeding in specific conditions but failing to generalize. ATLAS with Adaptive-OPRO delivers stable performance across tested regimes, with certain model pairings achieving positive returns even in volatile and declining market conditions where most baseline strategies struggle. The order-level action space reveals distinct trading patterns across model families and supports attribution from analytical reasoning to execution behavior.
@@ 5.1 Optimization in Sequential Decision-Making
Return and risk-adjusted metrics demonstrate successful adaptation to market feedback. Models paired with Adaptive-OPRO achieve improved returns while maintaining or reducing drawdowns, with Sharpe ratio gains indicating that performance improvements stem from genuine strategic enhancement rather than increased risk-taking. For instance, in the volatile bearish regime (Table 1), GPT-o3 and GPT-o4-mini shift from negative baseline returns to substantial positive performance under Adaptive-OPRO, while Qwen3-235B moves from losses to positive returns. This pattern appears across diverse regimes, from volatile declining markets to range-bound and bullish conditions, suggesting that the optimization mechanism captures regime-appropriate behaviors rather than overfitting to specific conditions.
Win rate improvements reveal an impact on decision consistency. Adaptive-OPRO configurations generally achieve higher win rates alongside better returns, indicating more reliable decision-making rather than occasional large gains masking frequent losses. This consistency appears even in models that struggle under fixed prompts, suggesting that prompt evolution helps models discover and maintain effective trading patterns.
The reflection paradox emerges across models and regimes: reflection-based approaches not only fail to match Adaptive-OPRO but frequently underperform baseline prompts. This finding challenges assumptions about the universal value of reflective reasoning in sequential decision systems. When base prompts are already well-engineered, adding reflection mechanisms appears to introduce complexity without corresponding benefits, and in some cases actively degrades performance by disrupting effective decision patterns.
@@
LLY No News 4.07± 0.72 0.056± 0.016 7.84± 3.15 53.51± 6.67 25.33± 4.51
(Bearish/Volatile No Market Data -5.75± 0.76 -0.094± 0.017 11.32± 2.63 37.52± 4.87 18.33± 3.06
Regime) No News & No Market -6.86± 1.68 -0.078± 0.036 14.54± 3.30 43.94± 6.94 22.33± 1.15
@@
XOM No News -8.20± 1.64 -0.264± 0.069 9.09± 2.99 22.82± 13.65 35.00± 12.29
(Sideways No Market Data 0.01± 0.92 -0.011± 0.021 6.56± 1.58 46.55± 23.15 13.33± 3.06
Regime) No News & No Market -4.60± 0.70 -0.136± 0.026 7.01± 2.29 35.26± 13.09 21.00± 4.58
@@
NVDA No News 6.62± 0.25 0.090± 0.008 6.67± 0.36 41.96± 5.21 28.33± 4.62
(Bullish No Market Data 11.78± 1.76 0.216± 0.024 3.70± 0.86 70.24± 14.03 20.00± 5.57
Regime) No News & No Market 7.34± 2.79 0.110± 0.012 5.76± 2.01 63.84± 9.39 20.67± 1.53
@@
5.2 Agent Contribution Analysis
@@
Market Analyst is a core component across market regimes. Its removal consistently results in the most significant performance degradation, especially in challenging conditions such as the bearish regime, where technical context is crucial for decision-making. In the sideways regime, the absence of market analysis not only reduces returns but also lowers trading frequency, suggesting that agents lose confidence to act without a solid technical foundation. Notably, in bullish markets, ROI slightly improves when market data are excluded, suggesting that in up-trending markets social consensus and market news may offer cleaner entry signals.
@@
News analyst contributes regime-specific strategic value. In the bullish regime, news removal leads to lower returns as agents become more conservative, missing chances to capitalize on positive momentum. The sideways regime shows news analysis as critical, with its removal producing severe degradation-suggesting that sentiment analysis is essential when technical signals are ambiguous.
@@
Combination of News & Market Analyst offers insights into their interdependent value: across all regimes, removing both agents leads to substantial performance degradation, indicating that news and market signals offer complementary, non-redundant information. In the bearish regime, performance drops significantly, reflecting the importance of sentiment and technical context under volatility. In the sideways regime, the absence of both leads to unstable and unprofitable behavior. Even in the bullish regime, where market data alone may be less essential, combined removal clearly harms performance. These results suggest that each component contributes differently across regimes, with the combined removal producing regime-specific effects that differ from simple additive impacts.
5.3 Trading Behavior Across LLMs
The order-level action space reveals systematic behavioral differences across model families, with performance correlating broadly with general model capabilities. Examining variance across runs provides additional insight into decision stability and reliability.
GPT-o3 demonstrates sophisticated strategy formulation, systematically integrating inputs from specialized agents into coherent trading decisions. The model exhibits conservative risk management, which limits gains in strongly trending markets but enables consistent performance across regimes. Its prompt optimization shows clear incremental learning and strategic adaptation to feedback. Notably, GPT-o3 maintains relatively low variance across runs, indicating stable decision patterns.
GPT-o4-mini favors short-term risk control through frequent stop-losses and early profit-taking. This approach proves effective in volatile conditions but struggles to capture sustained trends. The model tends toward higher trading frequencies in certain regimes, though its prompt optimization capabilities remain strong and its performance variance stays moderate.
Qwen models show divergent behavior based on scale. Qwen3-235B exhibits selective trading with moderate frequencies and achieves positive returns in multiple regimes, suggesting effective risk-reward balancing. Qwen3-32B displays higher activity levels with more variable outcomes, indicating less stable decision patterns. Both models benefit substantially from Adaptive-OPRO, with the larger variant showing particularly strong optimization responses.
LLaMA 3.3-70B adopts simpler trading strategies with limited risk management sophistication. The model shows delayed responses to market shifts and occasional abrupt strategy changes. Interestingly, this straightforward approach performs well in the bullish regime, effectively capturing upward trends without overcomplicating execution, though it struggles in more nuanced conditions.
Claude Sonnet 4 exhibits distinct behavior depending on reasoning mode. Variance patterns across runs reveal important differences in reliability: certain configurations show substantially higher standard deviations, indicating less predictable decision-making. With extended thinking enabled, the model produces detailed analysis but shows mixed translation to execution quality. Without thinking, decision patterns become even more erratic. Both configurations demonstrate challenges in maintaining consistent performance across regimes.
These behavioral patterns emerge clearly through order-level specifications, where we can separate analytical quality from execution choices. Models that struggle often show sound market analysis but poor translation into position sizing or timing, while successful configurations demonstrate coherent integration from analysis through execution. The variance patterns further distinguish models with stable, reproducible strategies from those with erratic decision-making that undermines reliability.
5.4 LLM Optimization Capabilities
A key advantage of Adaptive-OPRO is that the optimization process produces interpretable instruction updates that can be evaluated from two perspectives: whether the optimized prompt aligns with the trading objective, and whether the model successfully follows it. This dual lens reveals distinct capabilities across model families.
GPT models produce consistent, well-structured prompt refinements that demonstrate clear understanding of optimization objectives. The evolved instructions show coherent strategic adjustments that directly address observed weaknesses, and the models reliably follow their own optimized prompts.
Qwen models demonstrate strong prompt optimization capabilities across both variants. The generated refinements show analytical depth in identifying performance patterns and proposing targeted improvements, with the larger Qwen3-235B exhibiting particularly sophisticated reasoning about instruction design.
LLaMA frequently hallucinates edits, reporting modifications that do not appear in the actual prompt text or proposing changes that contradict the optimization objective. This disconnect between claimed improvements and actual instruction updates undermines the reliability of its optimization process.
Claude models tend toward overly rigid, procedural instructions that limit adaptability. The optimized prompts often become increasingly prescriptive rather than strategically focused, potentially constraining the model’s ability to respond flexibly to changing market conditions.
Reflection mechanisms, while conceptually appealing, introduce noise in practice. Even capable models can become paralyzed by over-analysis, while less sophisticated models generate vague reflections or confidently misinterpret market signals. This finding suggests that meta-reasoning about past decisions may complicate rather than clarify sequential decision-making when base prompts are already well-tuned. Further examples and analysis of these behaviors are provided in Appendix G.
@@ 6 Conclusion
We introduced ATLAS, an LLM-based trading framework integrating Adaptive-OPRO for prompt optimization under delayed, noisy feedback with structured analyst inputs and an order-level decision interface linking reasoning to execution. Across regime-specific equity studies and multiple model families, Adaptive-OPRO improves over strong fixed prompts, while reflection-based tuning yields no stable gains. The order-level interface exposes model-specific trading behaviors and separates analytical quality from execution choices, yielding clearer attribution and more interpretable decision traces. ATLAS illustrates a practical paradigm for sequential decision-making with LLMs, combining data-driven prompt evolution, structured inputs, and execution-aware outputs to achieve reliable, auditable, and accountable trading performance.
@@
Our scope-controlled experiments reveal how prompt optimization interacts with sequential decisions but necessarily limit breadth. We study three equities over regime-specific, two-month windows with daily decisions; sufficient to isolate adaptation effects but not to generalize across assets or market conditions. Results should thus be read as behavioral evidence of Adaptive-OPRO, not as market-wide performance claims.
Agents operate in an order-level simulator that enforces trading semantics while abstracting market microstructure. This improves control and attribution but omits execution frictions, so absolute returns may differ under richer dynamics. We use end-of-day decisions with a windowed ROI score chosen for stability rather than universality across horizons or volatility regimes.
Each configuration runs three times due to resource limits, capturing stochastic variance but limiting statistical power. Comparisons isolate adaptation under a shared interface rather than across full system architectures. Order-level actions aid interpretability and error attribution. We do not provide a causal comparison to directional-only outputs. Results span several model families (GPT, Claude, LLaMA, Qwen), but behaviors may vary with architecture, scale, or decoding choices not covered here.
@@ Appendix B Technical Indicators Used in Market Analysis
The Market Analyst consumes OHLCV, volume, and session VWAP series from Polygon.io [^2] https://polygon.io for the specified instrument and evaluation window. Bars are retrieved at daily resolution and aligned to official U.S. market sessions, with corporate actions (splits and dividends) from Polygon used to adjust prices consistently with StockSim. All technical indicators described in this appendix are computed inside StockSim from these Polygon-derived bars. Days with incomplete or missing bars are excluded rather than backfilled, and no survivorship or lookahead adjustments are applied beyond standard split and dividend handling.
@@ C.1 News Analyst
The News Analyst distills market-relevant information from financial news streams for a given ticker. Inputs are retrieved from the Polygon.io API [^3] https://polygon.io as batches of timestamped items containing title, URL, summary, and keywords. The component produces a structured analysis along four dimensions that are stable across models and assets: Sentiment Assessment, Key Developments, Market Relevance, and Source Analysis. When headline-only context is insufficient, the analyst can fetch the full article text through an internal fetcher to improve coverage and reduce headline bias. The output is designed to be compact, auditible, and directly consumable by the Central Trading Agent; it does not generate trading signals.
@@ C.2 Fundamental Analyst
The Fundamental Analyst extracts trading-relevant structure from periodic corporate disclosures (earnings releases, financial statements) and corporate actions (dividends, splits). It runs at low frequency to mirror real reporting cadence, typically activating once or twice per evaluation window. Inputs are retrieved via Polygon.io [^4] https://polygon.io and normalized to a compact schema consumed by the Central Trading Agent. The module does not emit buy/sell signals; it summarizes material changes and likely catalysts.
@@ Appendix D Experiments
D.1 Evaluation Scope

### v2

@@ 6 Results
Tables 1 and 2 present the results of our experimental design evaluating ATLAS across diverse market conditions. The results show that Adaptive-OPRO consistently improves upon fixed prompts across models and market conditions, while reflection often deteriorates performance or provides inconsistent value. Non-LLM strategies demonstrate regime-dependent performance, with different technical approaches succeeding in specific conditions but failing to generalize. ATLAS with Adaptive-OPRO delivers stable performance across tested regimes, with certain model pairings achieving positive returns even in volatile and declining market conditions where most baseline strategies struggle. The order-level action space reveals distinct patterns across model families and supports attribution from analytical reasoning to execution behavior.
@@ 6.1 Optimization in Sequential Decision-Making
Return, risk-adjusted, and win-rate metrics jointly indicate successful adaptation to market feedback. Models paired with Adaptive-OPRO achieve higher returns while maintaining or reducing drawdowns, with Sharpe ratio gains showing that improvements arise from strategic enhancement rather than increased risk-taking. Crucially, these return gains are accompanied by higher win rates, indicating more consistent decision-making rather than sporadic large profits masking frequent losses. For example, in the volatile bearish regime (Table 1), GPT-o3 and GPT-o4-mini shift from negative baseline returns to strong positive performance under Adaptive-OPRO, while Qwen3-235B moves from losses to gains. This pattern persists across range-bound and bullish conditions, suggesting that prompt optimization captures regime-appropriate behavior rather than overfitting to specific market settings.
Comparisons to baseline performance. We examine how baseline decision quality relates to the gains from Adaptive-OPRO under volatile, declining markets. Baseline and Adaptive-OPRO ROI are moderately correlated ($r=0.64$), suggesting that stronger baselines maintain higher absolute returns after adaptation. However, the improvement over baseline shows no meaningful correlation ($r=0.05$) and an almost flat gradient ($\beta\approx 0.06$). This indicates that Adaptive-OPRO does not simply amplify existing strengths, but delivers improvements largely independent of initial performance. Similar trends hold for risk-adjusted metrics, implying that Adaptive-OPRO mainly alters decision behavior rather than scaling baseline profitability.
The reflection paradox. In contrast, reflection-based prompting Li et al. (2024) exhibits a markedly different behavior. In the volatile bearish regime, the improvement in ROI under reflection shows a strong negative correlation with baseline performance (r=0.78,p<0.05r=-0.78,\,p<0.05), accompanied by a pronounced negative performance gradient ($\beta=-0.61$), indicating that models with stronger baseline decision quality tend to deteriorate more when reflection is introduced. This suggests that reflection does not merely fail to improve performance, but can actively disrupt effective decision policies in high-noise environments. Rather than stabilizing behavior, reflection appears to amplify stochasticity and override useful heuristics, particularly for models that already exhibit competent baseline trading strategies, consistent with overthinking induced by redundant information. Further examples and analysis in App. H.
@@
Volatile Regime No News 4.07± 0.72 0.056± 0.016 7.84± 3.15 53.51± 6.67 25.33± 4.51
No Market Data -5.75± 0.76 -0.094± 0.017 11.32± 2.63 37.52± 4.87 18.33± 3.06
No News & No Market -6.86± 1.68 -0.078± 0.036 14.54± 3.30 43.94± 6.94 22.33± 1.15
@@
Sideways Regime No News -8.20± 1.64 -0.264± 0.069 9.09± 2.99 22.82± 13.65 35.00± 12.29
No Market Data 0.01± 0.92 -0.011± 0.021 6.56± 1.58 46.55± 23.15 13.33± 3.06
No News & No Market -4.60± 0.70 -0.136± 0.026 7.01± 2.29 35.26± 13.09 21.00± 4.58
@@
Bullish Regime No News 6.62± 0.25 0.090± 0.008 6.67± 0.36 41.96± 5.21 28.33± 4.62
No Market Data 11.78± 1.76 0.216± 0.024 3.70± 0.86 70.24± 14.03 20.00± 5.57
No News & No Market 7.34± 2.79 0.110± 0.012 5.76± 2.01 63.84± 9.39 20.67± 1.53
@@
6.2 Trading Behavior Across LLMs
The order-level action space reveals systematic behavioral differences across model families, with performance broadly correlating with general model capabilities. Beyond averages, variance across runs captures decision reliability, especially when timing and sizing errors are amplified.
GPT models exhibit distinct trading styles and adaptation patterns. GPT-o3 integrates inputs from specialized agents into coherent decisions, showing conservative risk management that can cap gains in strongly trending markets but delivers consistent performance across regimes. This manifests as robust returns with comparatively low drawdowns and low run-to-run variance, indicating stable execution. GPT-o4-mini emphasizes short-term risk control through frequent stop-losses and early profit-taking. This behavior aligns with stronger outcomes in volatile settings and more muted trend capture in sustained moves; it also tends toward higher trading frequency in some regimes. Still, Adaptive-OPRO generally improves its consistency and profitability relative to fixed prompting, with moderate variance suggesting a more reactive but still controlled policy.
Qwen models show divergent behavior based on scale. Qwen3-235B trades more selectively and, across several regimes, achieves stable positive outcomes under Adaptive-OPRO. Both tables reflect that prompt adaptation is important here: it often turns otherwise marginal/negative behavior into positive returns while keeping activity relatively restrained, consistent with risk-reward balancing. Qwen3-32B is more active and variable, with larger swings across runs and regimes. Adaptive-OPRO improves its behavior, typically reducing losses in adverse settings and strengthening performance in favorable ones, but residual variance suggests less stable execution than the larger variant.
LLaMA 3.3-70B adopts simpler trading strategies with limited risk-management sophistication. Qualitatively, it shows delayed responses to market shifts and occasional abrupt changes in stance, which correspond to weaker performance in more adversarial regimes. Interestingly, this straightforward behavior performs well in the bullish regime in our results, consistent with capturing upward drift without overcomplicating execution.
Claude Sonnet 4 varies depending on reasoning mode, with variance patterns revealing meaningful differences in reliability. Certain configurations exhibit markedly higher run-to-run variability, indicating less predictable decision-making. With extended thinking enabled, the model often produces detailed analysis but the results show mixed execution quality; without thinking, decisions become more erratic and consistency across regimes degrades, suggesting that the bottleneck is both analysis depth and subsequent order construction.
Overall, the key insight enabled by order-level specifications is that weaker configurations often generate plausible market analysis but fail in position sizing, timing, or order selection, whereas successful configurations consistently translate analysis into coherent execution.
6.3 LLM Optimization Capabilities
A key advantage of Adaptive-OPRO is that optimization yields interpretable instruction updates that we can assess along two axes: (i) whether the revised prompt is objectively aligned with the trading goal (e.g., explicit risk controls, sizing discipline, and when to trade), and (ii) whether those instructions are reflected in subsequent order-level behavior (frequency, timing, and position sizing). After manual inspection of the results, we observe clear family-level patterns. GPT models consistently produce well-structured, objective-aligned refinements that translate observed weaknesses into actionable constraints, and these updates tend to be followed in execution, consistent with their lower run-to-run variance. Qwen models also generate targeted improvements, with the larger Qwen3-235B producing more coherent and internally consistent instruction revisions, which aligns with its more stable selective trading behavior. In contrast, LLaMA often reports edits that are not present in the actual prompt or proposes changes that conflict with the stated objective, weakening the connection between optimization output and downstream execution. Claude models frequently shift toward increasingly procedural and restrictive prompts, which can reduce adaptability; notably, this prescriptiveness does not reliably translate into stable execution, as reflected by higher variance in several configurations. Examples of the observed patterns are provided in Appendix G.
6.4 ATLAS Ablation Study
@@
Market Analyst is a core component across market regimes. Its removal consistently results in the most significant performance degradation, especially in challenging conditions such as the bearish regime, where technical context is crucial for decision-making. In the sideways regime, the absence of market analysis not only reduces returns but also lowers trading frequency, suggesting that agents lose confidence to act without a solid technical foundation. Notably, in bullish markets, ROI slightly improves when market data is excluded, suggesting that in up-trending markets social consensus may offer cleaner entry signals.
@@
News analyst contributes regime-specific strategic value. In the bullish regime, news removal leads to lower returns as agents become more conservative. The sideways regime shows news analysis as critical, with its removal producing severe degradation, suggesting that sentiment analysis is essential when technical signals are ambiguous.
@@
Combination of News & Market Analyst highlights the complementary value of these signals. Across all regimes, removing both agents substantially degrades performance, showing that news and market data provide non-redundant information. In the bearish regime, the drop reflects the importance of sentiment and technical context under volatility, while in the sideways regime their absence produces unstable, unprofitable behavior. Even in bullish markets, combined removal harms performance, indicating that each component contributes differently across regimes and that their joint effect is not simply additive.
@@ 7 Conclusion
In this work, we introduce ATLAS, an LLM-based trading framework that combines Adaptive-OPRO for prompt optimization under delayed, noisy feedback with structured analyst inputs and an order-level interface. Across regimes and model families, Adaptive-OPRO outperforms tuned static prompts, while standard reflection proves inconsistent. The order-level interface reveals model-specific trading behaviors and separates analytical quality from execution choices, enabling clearer attribution and interpretability. ATLAS with Adaptive-OPRO provides a practical, reliable, auditable paradigm for sequential LLM decision-making.
@@
Following prior LLM-agent and market-simulation work, we focus on three liquid equities over two-month, regime-specific windows with daily decisions to reduce confounding from asset heterogeneity and shifting market structure. This isolates adaptation effects under a shared interface but does not support generalization across assets, sectors, horizons, or macro conditions. Results should be read as behavioral evidence about Adaptive-OPRO, not as market-wide performance claims.
Agents operate in an order-level simulator that enforces trading semantics while abstracting market microstructure: slippage, partial fills, latency, and intraday dynamics are not modeled. This prioritizes experimental control and error attribution, consistent with prior simulation-based evaluations, but absolute returns may differ under real execution frictions. End-of-day decisions provide stable feedback for optimization under delayed, noisy outcomes, but prevent agents from reacting to intraday moves or capturing timing-dependent behaviors.
Each configuration runs three times due to resource constraints, capturing stochastic variance but limiting statistical power. Comparisons isolate prompt adaptation under a shared order-level interface rather than varying full system architectures. While order-level actions improve interpretability by separating analysis from execution, we do not include a directional-only ablation for direct causal comparison. Finally, although we cover multiple model families (GPT, Claude, LLaMA, Qwen), behaviors may vary with architectures, scales, and training procedures beyond those studied here.
@@ Appendix B Technical Indicators Used in Market Analysis
The Market Analyst consumes OHLCV, volume, and session VWAP series from Massive [^2] https://massive.com for the specified instrument and evaluation window. Bars are retrieved at daily resolution and aligned to official U.S. market sessions, with corporate actions (splits and dividends) from Massive used to adjust prices consistently with StockSim. All technical indicators described in this appendix are computed inside StockSim from these Massive-derived bars. Days with incomplete or missing bars are excluded rather than backfilled, and no survivorship or lookahead adjustments are applied beyond standard split and dividend handling.
@@ C.1 News Analyst
The News Analyst distills market-relevant information from financial news streams for a given ticker. Inputs are retrieved from the Massive API [^3] https://massive.com as batches of timestamped items containing title, URL, summary, and keywords. The component produces a structured analysis along four dimensions that are stable across models and assets: Sentiment Assessment, Key Developments, Market Relevance, and Source Analysis. When headline-only context is insufficient, the analyst can fetch the full article text through an internal fetcher to improve coverage and reduce headline bias. The output is designed to be compact, auditible, and directly consumable by the Central Trading Agent; it does not generate trading signals.
@@ C.2 Fundamental Analyst
The Fundamental Analyst extracts trading-relevant structure from periodic corporate disclosures (earnings releases, financial statements) and corporate actions (dividends, splits). It runs at low frequency to mirror real reporting cadence, typically activating once or twice per evaluation window. Inputs are retrieved via Massive [^4] https://massive.com and normalized to a compact schema consumed by the Central Trading Agent. The module does not emit buy/sell signals; it summarizes material changes and likely catalysts.
@@ Appendix D Experiments
D.1 Experimental Setup
Market regimes in our evaluation are instantiated using highly liquid, publicly traded equities selected prior to experimentation based on transparent criteria. Specifically, assets are required to exhibit stable liquidity conditions, clearly identifiable regime-consistent price dynamics over the evaluation window, and minimal microstructure distortions. This ensures that observed agent behavior reflects regime characteristics rather than artifacts of illiquidity or asset-specific noise. Asset instantiations are chosen independently of model performance and without outcome-driven adjustment, with selection criteria emphasizing representativeness of regime dynamics and sectoral diversity to reduce the likelihood that results are driven by idiosyncratic company- or industry-level effects.
Concretely, the bearish-volatile regime is instantiated using Eli Lilly and Company (LLY), the sideways regime using Exxon Mobil Corporation (XOM), and the bullish regime using NVIDIA Corporation (NVDA). All assets are evaluated over the same fixed two-month window (Apr 28–Jun 28, 2025) with a daily decision interval, ensuring consistency in sequential decision-making across regimes.
Importantly, ATLAS is asset- and regime-agnostic by design: no asset-specific features or regime-dependent assumptions are encoded in the framework, and the same experimental protocol can be directly applied to alternative equities, broader asset sets, or different evaluation horizons without modification.
D.2 Evaluation Scope